Find an equation of the line.
Through (-3,-6); parallel to 4x + 3y=7

Answer by Alan3354(69443)

Find an equation of the line.
Through (-3,-6); parallel to 4x + 3y=7
---------
Find the slope of 4x + 3y = 7
Solve for y
--> y = (-4/3)x + 7/3
That's the slope-intercept form. The slope, m, is the coefficient of x
m = -4/3
-----
Parallel lines have the same slope.
Use y-y1 = m*(x-x1) where (x1,y1) is the point (-3,-6)
----
y+6 = (-4/3)*(x+3) is an equation that fits.